# Fine-grained Image-to-LiDAR Contrastive Distillation with Visual Foundation Models

### v1

@@ Abstract
Contrastive image-to-LiDAR knowledge transfer, commonly used for learning 3D representations with synchronized images and point clouds, often faces a self-conflict dilemma. This issue arises as contrastive losses unintentionally dissociate features of unmatched points and pixels that share semantic labels, compromising the integrity of learned representations. To overcome this, we harness Visual Foundation Models (VFMs), which have revolutionized the acquisition of pixel-level semantics, to enhance 3D representation learning. Specifically, we utilize off-the-shelf VFMs to generate semantic labels for weakly-supervised pixel-to-point contrastive distillation. Additionally, we employ von Mises-Fisher distributions to structure the feature space, ensuring semantic embeddings within the same class remain consistent across varying inputs. Furthermore, we adapt sampling probabilities of points to address imbalances in spatial distribution and category frequency, promoting comprehensive and balanced learning. Extensive experiments demonstrate that our approach mitigates the challenges posed by traditional methods and consistently surpasses existing image-to-LiDAR contrastive distillation methods in downstream tasks. The source code is available at https://github.com/Eaphan/OLIVINE.
@@ 1 Introduction
LiDAR sensors deliver critical information in the 3D environment, crucial for applications such as autonomous driving [66; 63]. State-of-the-art neural networks have shown promising performance on point clouds processing, which rely on extensive annotated datasets [25; 65; 67]. However, the process of annotating point clouds is not only time-consuming but also costly, presenting significant challenges in terms of scalability and practicality [31]. Self-supervision offers a solution by leveraging vast quantities of unlabeled data to pre-train networks and subsequently fine-tuning them with a smaller, labeled dataset. This approach significantly reduces the reliance on extensive annotated data sets [9].
@@
Unfortunately, existing contrastive distillation methods are hindered by several critical limitations. Firstly, a "self-conflict" issue arises during pre-training, where (super)pixels that belong to the same category as an anchor (super)point but do not directly correspond are simply treated as negative samples (see Fig.1(a)). This approach neglects the inherent semantic connections within the same category, leading to conflicts in the learning process where beneficial relationships might be disregarded. This problem is magnified by the contrastive loss’s intrinsic hardness-aware characteristic, which results in the most substantial gradient influences derived from negative samples that are semantically the most similar [50; 36]. While ST-SLidR [36] has introduced a semantically tolerant loss to mitigate this issue, the absence of a robust high-level semantic understanding could not fundamentally change
@@ 2 Related Work
3D Representation Learning. Recent advancements in 3D self-supervised learning have closely paralleled innovations in the image domain, extending these methods to diverse 3D contexts such as object-level point clouds [42; 51; 21], indoor scenes [52; 11; 69; 57; 8; 27; 58; 49], and outdoor settings [29; 3; 60; 37; 16]. These techniques are grounded in contrastive learning [55; 60; 37], mask modeling [61], and other pretext tasks [3; 64]. PPKT [34] utilized the InfoNCE loss to facilitate the 3D network in distilling rich knowledge from the 2D image backbone. Sautier et al.[43] pioneered a superpixelto-superpoint contrastive loss for self-supervised 2D-to-3D representation distillation. Building on this, Mahmoud et al. [36] enhanced the approach by incorporating a semantically tolerant contrastive constraint and a class-balancing loss. Liu et al.[33] further refined these techniques through semantic-aware spatial and temporal consistency regularization, advancing feature learning. Moreover, Zhang et al. [62] explored cross-modality contrastive distillation across not only the image plane but also the bird-eye views.
Visual Foundation Models. The advent of powerful visual neural networks, trained on extensive datasets [39; 24] or through cutting-edge self-supervised learning techniques [6; 10; 19], catalyzed significant advancements within the community. Notably, the Segment Anything Model (SAM) [24] initiated a new paradigm in general-purpose image segmentation, demonstrating remarkable zero-shot transfer capabilities across a variety of downstream tasks. Building on this, Grounded-SAM [40] enhanced the model by incorporating elements from Grounding-DINO [32], an open-set object detector capable of recognizing and classifying previously unseen objects during training [32]. In our work, we leverage the inherent semantic awareness of these VFMs to generate weak semantic labels, which are crucial for our supervised contrastive distillation framework.
3D Scene Understanding. Traditional approaches to 3D scene understanding primarily utilize paradigms based on raw points [47; 14; 7; 53], voxels [15; 12], range views [25; 48], and multi-view fusion [17; 56]. Despite the effectiveness of these methods in capturing detailed environmental features, they are significantly constrained by their dependence on large-scale annotated data. Acquiring and annotating this data is both time-consuming and costly, limiting the scalability of 3D segmentation models [31]. To reduce the reliance on large annotated datasets, recent studies have also turned to semi-supervised [26; 20], weakly-supervised [30; 13], and active learning techniques [35; 54].
@@ 3.1 Baseline Architecture
We follow the existing work [34] to perform the basic point-to-pixel contrastive distillation, based on which we build our whole pipeline. Starting with point cloud and image inputs, we employ distinct encoders for feature extraction. The 3D features are extracted using an encoder $f_{\mathrm{3D}}(\cdot):\mathbb{R}^{N\times 3}\to\mathbb{R}^{N\times C_{\mathrm{3D}}}$, which processes the point clouds to produce features of dimension $C$ per point. For image features, we use an encoder $f_{2D}(\cdot):\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{H^{\prime}\times W^{\prime}\times C_{\mathrm{2D}}}$, initialized with weights from pre-trained image models. This setup facilitates knowledge transfer from the 2D domain to the 3D domain through contrastive learning. For the computation of contrastive loss, we design trainable projection heads, $h_{\mathrm{2D}}^{\mathrm{pp}}$ for 2D features and $h_{\mathrm{3D}}^{\mathrm{pp}}$ for 3D features, both aligning the features into a unified dimensional space, $C$. Specifically, the 3D projection head $h_{\mathrm{3D}}^{\mathrm{pp}}$ is a linear layer with $\ell_{2}$-normalization, converting 3D features to a normalized $C$-dimensional space. Similarly, the 2D projection head $h_{\mathrm{2D}}^{\mathrm{pp}}$, a convolution layer with a 1$\times$1 kernel followed by a bi-linear interpolation layer adjusting the spatial dimensions by a factor of 4, also applies $\ell_{2}$-normalization.
@@ 3.2 Weakly-supervised Contrastive Distillation
Existing methods often mistakenly treat unmatched points and pixels that share semantic labels as negative pairs. This practice overlooks the intrinsic semantic connections within the same categories, leading to potential conflicts in the learning process where beneficial relationships are ignored. To address this, we utilize the Segment Anything Model, which adeptly interprets and translates semantic cues from text prompts into precise semantic segmentation of images. This application of SAM allows us to generate high-quality semantic labels without repetitive training, enhancing the learning process. We represent these labels as $Y^{\mathrm{co}}=\{y^{\mathrm{co}}_{i}\}_{i=1}^{M}$, where each label corresponds to a specific point-pixel pair.
@@ 3.3 Semantic-guided Consistency Regularization
Parameter Updating. Specifically, we conduct the semantic-guided consistency regularization in an two-stage framework. First, we update the parameter of $\mathrm{vMF}(z;\mu_{k},\kappa_{k}))$ with the features $\bm{G}^{\mathrm{3D}}$ extracted from point cloud branch. During training, we obtain the statistical value of feature embeddings via the EMA (Exponential Moving Average) algorithm, following:
@@
$\delta(z-\bm{G}^{\mathrm{3D}}_{i})=\left\{\begin{array}[]{l l}{0}&{z\not=\bm{G}^{\mathrm{3D}}_{i}}\\ {\infty}&{z=\bm{G}^{\mathrm{3D}}_{i}}\end{array}\right.$ (12)
@@ 4 Experiments
4.1 Experimental Setup of 3D Pretraining
Pre-training Details. We utilize momentum SGD for optimization, setting the initial learning rate at 0.5 and 0.01 for SR-UNet34 and VoxelNet respectively, with a momentum of 0.9 and a weight decay of 1e-4. To adjust the learning rate, we employ a cosine annealing scheduler [4] that gradually reduces it from the initial value to 0 over 50 epochs. The 3D network is pre-trained for these 50 epochs on four NVIDIA-3090 GPUs, processing a total batch size of 16, unless specified otherwise. For data augmentation, we incorporate several techniques. For the point cloud data, we apply random rotations around the z-axis, randomly flip the x and y axes, and omit points within a randomly selected cuboid, following the method described in [68]. For image data, augmentations include random horizontal flips and random crop-resize operations. In terms of generating weak semantic labels, the prompts we provided to the Grounded-SAM [40] encompass a total of 13 object categories: barrier, bicycle, bus, car, motorcycle, pedestrian, traffic cone, truck, road, sidewalk, terrain, vegetation, and building.
Table 1: Comparison of various pre-training techniques for semantic segmentation tasks using either finetuning or linear probing. This evaluation uses different proportions of accessible annotations from the nuScenes datasets and presents the mean Intersection over Union (mIoU) scores on the nuScenes validation set.
  Initialization Present at Lin. prob. Finutune
100% 1% 5% 10% 25% 100%
Random - 8.1 30.30 47.84 56.15 65.48 74.66
PointContrast [55] ECCV’20 21.90 32.50 - - - -
DepthContrast [68] ICCV’21 22.10 31.70 - - - -
PPKT [34] Arxiv’21 35.90 37.80 53.74 60.25 67.14 74.52
SLidR [43] CVPR’22 38.80 38.30 52.49 59.84 66.91 74.79
ST-SLidR [36] CVPR’23 40.48 40.75 54.69 60.75 67.70 75.14
HVDistill [62] IJCV’24 39.50 42.70 56.60 62.90 69.30 76.60
Seal [33] NeurIPS’23 44.95 45.84 55.64 62.97 68.41 75.60
Ours - 47.30 46.12 57.51 63.04 69.39 76.13
@@ 4.2 Transfer on Semantic Segmentation
Evaluation Protocol. We evaluate the learned representations for semantic segmentation on nuScenes-lidarseg and SemanticKITTI datasets. The nuScenes-lidarseg and SemanticKITTI datasets contain 16 and 19 semantic categories for validation, respectively. We modify the network by adding a 3D convolutional layer to the pre-trained backbone as the segmentation head. Basically, we finetune the whole network on different portions of annotated data. Following previous works [43; 36], we finetune the network for 100 epochs with a batch size of 10 and 16 on SemanticKITTI and nuScenes-lidarseg, respectively. The initial learning rate of the backbone and the segmentation head are set to 0.05 and 2.0, respectively. When utilizing 1% of the annotated data, the network is fine-tuned for 100 epochs, whereas for other percentages, it is fine-tuned for 50 epochs. In another protocol, we evaluate the quality of learned representation by linear probing. Different from the setting of finetuning, we optimize only the added segmentation head and keep the weights of backbone $f_{\mathrm{3D}}$ frozen on the nuScenes-lidarseg dataset. For both protocols, the training objective is a linear combination of the cross-entropy loss and the Lovász-Softmax loss [2].
@@
Table 2: Comparison of various pre-training techniques for semantic segmentation through finetuning on the SemanticKITTI dataset with only 1% annotated data.
  Initialization Random Point- Contrast Depth- Contrast PPKT SLidR ST-SLidR Seal Ours
mIoU (%) 39.50 41.10 41.50 44.00 44.60 44.72 46.63 47.35
Improvement - +0.60 +1.00 +4.50 +5.10 +5.22 +7.13 +7.85
@@
Ours 47.4 +7.9 56.5 +10.8 58.2 +6.7 59.5 +3.5 59.7 +2.8
Results of Linear Probing. Under the linear probing scenario, our method achieves the highest mIoU of 47.30%, surpassing the previous state-of-the-art method, Seal [33], which records a mIoU of 44.95% (see Table 1). This performance indicates a significant improvement in extracting useful features directly from pre-trained models without additional training of the 3D backbone.
Results of Fine-tuning. For the fine-tuning on nuScenes, our consistently excel, particularly at smaller data proportions. With only 1% of the training data of the nuScenes, our method achieves a mIoU of 46.12%, outperforming the HVDistill [62] that had a mIoU of 42.70%. This trend persists across other data proportions, with our method consistently leading or closely competing with the best results, particularly at 57.51% mIoU with 5% of the data and 63.04% mIoU with 10% of the data. The qualitative results are presented in Fig. 3.
As shown in Table 2, our method also achieves state-of-the-art performance on the SemanticKITTI dataset. With only 1% of the data annotated, our method achieved a mIoU of 47.35%, which is a substantial improvement of 7.85% over the baseline. We also fine-tuned various models on the SemanticKITTI dataset to assess their performance across a spectrum of annotated data availability, from as low as 1% to full data utilization (see Table 3). Specifically, with only 1% of the data annotated, our method achieved a mIoU of 47.35%, which is a substantial improvement of 7.85% over the baseline. This trend continues as more data becomes available; with 5% of data, our model further increased its lead by achieving a mIoU of 56.5%, reflecting an improvement of 10.8%.

### v3

@@ Abstract
Contrastive image-to-LiDAR knowledge transfer, commonly used for learning 3D representations with synchronized images and point clouds, often faces a self-conflict dilemma. This issue arises as contrastive losses unintentionally dissociate features of unmatched points and pixels that share semantic labels, compromising the integrity of learned representations. To overcome this, we harness Visual Foundation Models (VFMs), which have revolutionized the acquisition of pixel-level semantics, to enhance 3D representation learning. Specifically, we utilize off-the-shelf VFMs to generate semantic labels for weakly-supervised pixel-to-point contrastive distillation. Additionally, we employ von Mises-Fisher distributions to structure the feature space, ensuring semantic embeddings within the same class remain consistent across varying inputs. Furthermore, we adapt sampling probabilities of points to address imbalances in spatial distribution and category frequency, promoting comprehensive and balanced learning. Extensive experiments demonstrate that our approach mitigates the challenges posed by traditional methods and consistently surpasses existing image-to-LiDAR contrastive distillation methods in downstream tasks. Code is available at https://github.com/Eaphan/OLIVINE.
@@ 1 Introduction
LiDAR sensors deliver critical information in the 3D environment, crucial for applications such as autonomous driving [78; 75; 79]. State-of-the-art neural networks have shown promising performance on point clouds processing, which rely on extensive annotated datasets [29; 77; 17]. However, the process of annotating point clouds is not only time-consuming but also costly, presenting significant challenges in terms of scalability and practicality [36]. Self-supervision offers a solution by leveraging vast quantities of unlabeled data to pre-train networks and subsequently fine-tuning them with a smaller, labeled dataset. This approach significantly reduces the reliance on extensive annotated data sets [10].
@@
Unfortunately, existing contrastive distillation methods are hindered by several critical limitations. Firstly, a "self-conflict" issue arises during pre-training, where (super)pixels that belong to the same category as an anchor (super)point but do not directly correspond are simply treated as negative samples (see Fig.1(a)). This approach neglects the inherent semantic connections within the same category, leading to conflicts in the learning process where beneficial relationships might be disregarded. This problem is magnified by the contrastive loss’s intrinsic hardness-aware characteristic, which results in the most substantial gradient influences derived from negative samples that are semantically the most similar [59; 41]. While ST-SLidR [41] has introduced a semantically tolerant loss to mitigate this issue, the absence of a robust high-level semantic understanding could not fundamentally change
@@ 2 Related Work
3D Representation Learning. Recent advancements in 3D self-supervised learning have closely paralleled innovations in the image domain, extending these methods to diverse 3D contexts such as object-level point clouds [49; 60; 23], indoor scenes [61; 12; 81; 68; 9; 32; 69; 58], and outdoor settings [34; 4; 71; 42; 18]. These techniques are grounded in contrastive learning [66; 71; 42], mask modeling [72], and other pretext tasks [4; 76]. PPKT [39] utilized the InfoNCE loss to facilitate the 3D network in distilling rich knowledge from the 2D image backbone. Sautier et al.[50] pioneered a superpixelto-superpoint contrastive loss for self-supervised 2D-to-3D representation distillation. Building on this, Mahmoud et al. [41] enhanced the approach by incorporating a semantically tolerant contrastive constraint and a class-balancing loss. Liu et al.[38] further refined these techniques through semantic-aware spatial and temporal consistency regularization, advancing feature learning. Moreover, Zhang et al. [73] explored cross-modality contrastive distillation across not only the image plane but also the bird-eye views.
Visual Foundation Models. The advent of powerful visual neural networks, trained on extensive datasets [46; 27] or through cutting-edge self-supervised learning techniques [7; 11; 21], catalyzed significant advancements within the community. Notably, the Segment Anything Model (SAM) [27] initiated a new paradigm in general-purpose image segmentation, demonstrating remarkable zero-shot transfer capabilities across a variety of downstream tasks. Building on this, Grounded-SAM [47] enhanced the model by incorporating elements from Grounding-DINO [37], an open-set object detector capable of recognizing and classifying previously unseen objects during training [37]. In our work, we leverage the inherent semantic awareness of these VFMs [84] to generate weak semantic labels, which are crucial for our supervised contrastive distillation framework.
3D Scene Understanding. Traditional approaches to 3D scene understanding primarily utilize paradigms based on raw points [55; 16; 8; 62], voxels [17; 13], range views [29; 56], and multi-view fusion [19; 67]. Despite the effectiveness of these methods in capturing detailed environmental features, they are significantly constrained by their dependence on large-scale annotated data. Acquiring and annotating this data is both time-consuming and costly, limiting the scalability of 3D segmentation models [36]. To reduce the reliance on large annotated datasets, recent studies have also turned to semi-supervised [31; 22], weakly-supervised [35; 15], and active learning techniques [40; 65].
@@ 3.1 Baseline Architecture
We follow the existing work [39] to perform the basic point-to-pixel contrastive distillation, based on which we build our whole pipeline. Starting with point cloud and image inputs, we employ distinct encoders for feature extraction. The 3D features are extracted using an encoder $f_{\mathrm{3D}}(\cdot):\mathbb{R}^{N\times 3}\to\mathbb{R}^{N\times C_{\mathrm{3D}}}$, which processes the point clouds to produce features of dimension $C_{\mathrm{3D}}$ per point. For image features, we use an encoder $f_{2D}(\cdot):\mathbb{R}^{H\times W\times 3}\to\mathbb{R}^{H^{\prime}\times W^{\prime}\times C_{\mathrm{2D}}}$, initialized with weights from pre-trained image models. This setup facilitates knowledge transfer from the 2D domain to the 3D domain through contrastive learning. For the computation of contrastive loss, we design trainable projection heads, $h_{\mathrm{2D}}^{\mathrm{pp}}$ for 2D features and $h_{\mathrm{3D}}^{\mathrm{pp}}$ for 3D features, both aligning the features into a unified dimensional space. Specifically, the 3D projection head $h_{\mathrm{3D}}^{\mathrm{pp}}$ is a linear layer with $\ell_{2}$-normalization, converting 3D features to a normalized $C$-dimensional space. Similarly, the 2D projection head $h_{\mathrm{2D}}^{\mathrm{pp}}$, a convolution layer with a 1$\times$1 kernel followed by a bi-linear interpolation layer adjusting the spatial dimensions by a factor of 4, also applies $\ell_{2}$-normalization.
@@ 3.2 Weakly-supervised Contrastive Distillation
Existing methods [39; 50; 38] often directly treat unmatched points and pixels that share semantic labels as negative pairs. This practice overlooks the intrinsic semantic connections within the same categories, leading to potential conflicts in the learning process where beneficial relationships are ignored. To address this, we utilize the Segment Anything Model, which adeptly interprets and translates semantic cues from text prompts into precise semantic segmentation of images. This application of SAM allows us to generate high-quality semantic labels without repetitive training, enhancing the learning process. We represent these labels as $Y^{\mathrm{co}}=\{y^{\mathrm{co}}_{i}\}_{i=1}^{M}$, where each label corresponds to a specific point-pixel pair.
@@ 3.3 Semantic-guided Consistency Regularization
Parameter Updating. Specifically, we conduct the semantic-guided consistency regularization in a two-stage framework. First, we update the parameter of $\mathrm{vMF}(z;\mu_{k},\kappa_{k}))$ with the features $\bm{G}^{\mathrm{3D}}$ extracted from point cloud branch. During training, we obtain the statistical value of feature embeddings via the EMA (Exponential Moving Average) algorithm, following:
@@
$\delta(z-\bm{G}^{\mathrm{3D}}_{i})=\left\{\begin{array}[]{l l}{0},&{z\not=\bm{G}^{\mathrm{3D}}_{i}}\\ {\infty},&{z=\bm{G}^{\mathrm{3D}}_{i}}\end{array}\right.$ (12)
@@ 3.4 Density and Category-aware Sampling Strategy
Table 1: Comparison of various pre-training techniques for semantic segmentation tasks using either finetuning or linear probing (LP). This evaluation uses different proportions of accessible annotations from the nuScenes or SemanticKITTI datasets and presents the mean Intersection over Union (mIoU) scores on the validation set.
  Initialization Present at nuScenes KITTI
LP 1% 5% 10% 25% 100% 1%
Random - 8.1 30.30 47.84 56.15 65.48 74.66 39.50
PointContrast [66] ECCV’20 21.90 32.50 - - - - 41.10
DepthContrast [80] ICCV’21 22.10 31.70 - - - - 41.50
PPKT [39] Arxiv’21 35.90 37.80 53.74 60.25 67.14 74.52 44.00
SLidR [50] CVPR’22 38.80 38.30 52.49 59.84 66.91 74.79 44.60
ST-SLidR [41] CVPR’23 40.48 40.75 54.69 60.75 67.70 75.14 44.72
Seal [38] NeurIPS’23 44.95 45.84 55.64 62.97 68.41 75.60 46.63
Ours NeurIPS’24 50.09 50.58 60.19 65.01 70.13 76.54 49.38
@@ 4.1 Transfer on Semantic Segmentation
Evaluation Protocol. We evaluate the learned representations for semantic segmentation on nuScenes-lidarseg and SemanticKITTI datasets. The nuScenes-lidarseg and SemanticKITTI datasets contain 16 and 19 semantic categories for validation, respectively. We modify the network by adding a 3D convolutional layer to the pre-trained backbone as the segmentation head. Basically, we finetune the whole network on different portions of annotated data. Following previous works [50; 41], we finetune the network for 100 epochs with a batch size of 10 and 16 on SemanticKITTI and nuScenes-lidarseg, respectively. The initial learning rate of the backbone and the segmentation head are set to 0.05 and 2.0, respectively. When utilizing 1% of the annotated data, the network is fine-tuned for 100 epochs, whereas for other percentages, it is fine-tuned for 50 epochs. In another protocol, we evaluate the quality of learned representation by linear probing. Different from the setting of finetuning, we optimize only the added segmentation head and keep the weights of backbone $f_{\mathrm{3D}}$ frozen on the nuScenes-lidarseg dataset. For both protocols, the training objective is a linear combination of the cross-entropy loss and the Lovász-Softmax loss [3].
@@
Ours 49.4 +9.9 57.5 +11.8 60.3 +8.8 60.9 +4.9 61.1 +4.2
Results of Linear Probing. Under the linear probing scenario, our method achieves the highest mIoU of 50.09%, surpassing the previous state-of-the-art method, Seal [38], which records a mIoU of 44.95% (see Table 1). This performance indicates a significant improvement in extracting useful features directly from pre-trained models without additional training of the 3D backbone.
Results of Fine-tuning. For the fine-tuning on nuScenes, our consistently excel, particularly at smaller data proportions. With only 1% of the training data of the nuScenes, our method achieves a mIoU of 50.58%. This trend persists across other data proportions, with our method consistently leading or closely competing with the best results, particularly at 60.19% mIoU with 5% of the data and 65.01% mIoU with 10% of the data. The qualitative results are presented in Fig. 3.
We also fine-tuned various models on the SemanticKITTI dataset to assess their performance across a spectrum of annotated data availability, from as low as 1% to full data utilization (see Table 2). Besides, as shown in Table 3, our method consistently achieves state-of-the-art performance on another six datasets.
Table 3: Evaluation of various pretraining methods initially trained on the nuScenes dataset and subsequently fine-tuned on multiple downstream point cloud datasets. The mIoU scores are presented as percentages (%).
Method ScribbleKITTI RELLIS-3D SemanticPOSS SemanticSTF SynLiDAR DAPS-3D
1% 10% 1% 10% 50% 100% 50% 100% 1% 10% 50% 100%
Random 23.81 47.60 38.46 53.60 46.26 54.12 48.03 48.15 19.89 44.74 74.32 79.38
PPKT [39] 36.50 51.67 49.71 54.33 50.18 56.00 50.92 54.69 37.57 46.48 78.90 84.00
SLidR [50] 39.60 50.45 49.75 54.57 51.56 55.36 52.01 54.35 42.05 47.84 81.00 85.40
Seal [38] 40.64 52.77 51.09 55.03 53.26 56.89 53.46 55.36 43.58 49.26 81.88 85.90
Ours 44.91 54.96 53.47 58.21 55.70 58.51 56.65 60.42 46.34 52.78 83.63 86.84
@@ 4.2 Transfer on 3D Object Detection
Remark. Compared to the semantic segmentation task, the model architecture for object detection is more complex. Besides the 3D backbone, 3D detectors typically project features to a BEV plane, followed by a 2D convolutional network and RoI operations. These crucial components were not pre-trained, which may limit the overall performance gain from our pre-training approach. It’s important to note that semantic segmentation and object detection use different metrics and scales, making direct performance comparisons improper. The nature of these tasks and their evaluation criteria inherently lead to varying degrees of improvement when applying our proposed method.